Resolve that cos3A-cos5A by sin3A+sin5A = tanA


1
Expert's answer
2021-02-24T07:49:16-0500

"\frac{2sin(\frac{5A−3A}{2})cos(\frac{5A+3A}{2})}{2cos(\frac{5A+3A}{2})cos(\frac{5A−3A}{2})}"

"\frac{sinA}{cosA}"

"tanA"